Mendelian Genetics
- Background
- Mendel used pea plants because they were cheap, had variety, easy to grow, had a short generation type/many offspring/identifiable traits/easy to control pollination Crossed true-breeding plants (all individuals of line have = characteristics) w/ contrasting traits (purple and white) P = parental generation Crossed 2 plants in P to become results in F1 generation
Subjects Alleles results Blending Pea plants Blend No intermediate phenotypes (no light purple) and lost phenotypes reappear Particular inheritance Pea plants discrete Heritable traits are discrete units (genes) that determine characters, each char. Controlled by 2 factors (genes) from parents
Particular inheritance (3:1 inheritance pattern) Alleles (traits) are discrete 2 alleles from each parent Dominant and recessive alleles 2 laws of inheritance Alleles are alternate versions of gene (2n = 2 of each allele), each one @ same locus on homologous chromosomes, segment of DNA Homozygous: 2 alleles identical P generation Heterozygous: 2 alleles different P generation Dominant allele: determines appearance Recessive allele: has no effect on appearance Law of segregation: each gamete gets 1 of each allele, 2 copies of each gene (allele) segregate during meiosis, egg/sperm only gets 1 of 2 alleles; ANAPHASE Law of independent assortment: action of sorting, genes on different chromosomes randomly/independently assort during gamete formation bc of random orientation of tetrads in METAPHASE I, leads to gen. recombination/> variation Genetic crosses Punnett squares show mendel’s laws and possible babies Phenotype: PHYSICAL/expressed type Genotype: genetic type (YY, Yy, yy) Monohybrid cross: cross of 1 character/allele Dominant is uppercase, recessive is lowercase Each sperm/egg has 1 allele of all alleles parents have If DD and dd crossed: 1:2:1 genotypic ratio, 3:1 phenotypic ratio in F2 Test cross: cross unknown genotype with known genotype that is recessive Probability Multiplication rule: probability of independent events, 1 event doesn’t affect another P(This) and P(that) Addition rule: probability of mutually exclusive events, can’t occur at same time P(this) or P(that) Ex: mother is Bb and father is Bb, probability that b from mom and B from dad OR B from mom and b from dad? .5*.5=.25; .5*.5 = .25; .25 + .25 = .5